# Mamba Neural Operator: Who Wins? Transformers vs. State-Space Models for PDEs

### v2

@@ IV-C Why the Winner Wins: Breaking Down Mamba’s Win
Figure 7 shows the $\Delta$ log magnitudes across the normalised depth for the Galerkin attention, Softmax attention, and Mamba versions of GNOT. The Galerkin and Softmax versions exhibit sharp fluctuations, indicating instability and inconsistent feature extraction at different depths. In contrast, Mamba maintains a steady and flat profile, reflecting robust and stable feature extraction. The gray and white bands indicate the alternating roles of the operator and NLP components, further emphasising Mamba’s balanced performance across layers, making it ideal for handling complex PDEs.
TABLE III: Comparisons with different query positions using nRMSE.
Method Identical Diagonal
@@
TABLE IV: Comparisons with different dataset sizes using nRMSE.
Method 9K 5K 2K 1K
@@
IV-D Ablation Study: Final Battles, Winner Takes All
@@
Table III compares nRMSE performance across two query scenarios: Identical positions (input and query points are the same) and Diagonal positions (shifted inputs creating a mismatch). Experiments are done using Darcy Flow. While prior work shows that Transformers handle inconsistent input-query positions well[10], our results demonstrate a clear advantage of Mamba in both configurations. For Identical query positions, Mamba version achieves the lowest error, outperforming OFormer and its softmax variant, demonstrating Mamba’s superior ability to capture relationships when input and query points are perfectly aligned. For Diagonal query positions, where inputs and queries are misaligned, Mamba achieves the best performance compared to OFormer and its variant, demonstrating its superior ability to generalise under spatial shifts.
@@
Table IV compares nRMSE performance across different dataset sizes for GNOT, GNOT with softmax attention (w/ S.A.), and GNOT with Mamba. Experiments are carried out using Darcy Flow. As dataset size decreases, Mamba consistently achieves the lowest error, demonstrating superior performance and robustness in data-scarce scenarios. For instance, with the smallest dataset (1K), Mamba achieves an nRMSE of 0.0617, significantly lower than GNOT’s and GNOT w/ S.A.’s, showcasing its resilience and generalisation capability even with limited data. This highlights Mamba’s efficiency in learning meaningful representations with fewer data points, making it a powerful choice for real-world applications where data availability is a constraint.
V Conclusion
@@ Acknowledgments
CWC is supported by the Swiss National Science Foundation (SNSF) under grant number 20HW-1_220785. It also acknowledge CMI, University of Cambridge. JH was supported in part by the Imperial College Bioengineering Department PhD Scholarship and the UKRI Future Leaders Fellowship (MR/V023799/1). GY were supported in part by the ERC IMI (101005122), the H2020 (952172), the MRC (MC/ PC/21013), the Royal Society (IEC NSFC211235), the NVIDIA Academic Hardware Grant Program, the SABER project supported by Boehringer Ingelheim Ltd, Wellcome Leap Dynamic Resilience, and the UKRI Future Leaders Fellowship (MR/V023799/1). CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC advanced career fellowship EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards 215733/Z/19/Z and 221633/Z/20/Z, CCMI and the Alan Turing Institute. AIAR gratefully acknowledges the support from Yau Mathematical Sciences Center, Tsinghua University.

### v3

@@ 4.3 Why the Winner Wins: Breaking Down Mamba’s Win
Figure 9 shows the $\Delta$ log magnitudes across the normalised depth for the Galerkin attention, Softmax attention, and Mamba versions of GNOT. The Galerkin and Softmax versions exhibit sharp fluctuations, indicating instability and inconsistent feature extraction at different normalised depths. In contrast, Mamba maintains a steady and flat profile, reflecting robust and stable feature extraction. The gray and white bands indicate the alternating roles of the operator and NLP components, further emphasising Mamba’s balanced performance across layers, making it ideal for handling complex PDEs.
Table 3: Comparisons with different query positions
using nRMSE.
Method Query Positions
Identical Diagonal
@@
Table 4: Comparisons with different dataset sizes using nRMSE.
Method Dataset Sizes
9K 5K 2K 1K
@@
Table 5: Comparisons on query positions and dataset sizes using nRMSE.
Method (a) Query Positions (b) Dataset Sizes
Identical Diagonal 9K 5K 2K 1K
OFormer 0.0253 0.0318
w/S.A. 0.0324 0.0382 0.0394 0.0400 0.0526 0.0776
w/Mamba 0.0244 0.0314 0.0367 0.0376 0.0481 0.0617
GNOT 0.0485 0.0567 0.0777 0.1174
Table 6: Quantitative comparisons on Darcy Flow across three Transformer-based architectures with Galarkin attention, softmax attention, and Mamba Neural Operator (MNO). Metrics include FLOPs, parameter count, inference time, and GPU memory usage
DarcyFlow
Method Type FLOPs(G) Parameters (M) Inference time (ms) GPU Memory Usage (GIB)
GNOT (Hao et al, 2023) Galerkin 29.71 5.59 12.04 0.15
w/S.A. Softmax 235.26 5.59 162.93 8.12
w/Mamba (MNO) Mamba 45.36 7.65 19.70 0.48
G.T. (Cao et al, 2021) Galerkin 85.59 26.24 15.85 0.27
w/S.A. Softmax 503.94 26.24 132.58 18.67
w/Mamba (MNO) Mamba 116.99 33.45 32.56 0.78
OFormer (Li et al, 2023) Normalised 116.34 18.04 39.66 1.11
w/S.A. Softmax 977.14 19.10 127.64 4.83
w/Mamba (MNO) Mamba 60.13 10.89 22.89 1.13
4.4 Ablation Study: Final Battles, Winner Takes All
Figure 9: Relative L2 error on SW2D across time steps, comparing the performance of our Mamba Neural Operator (MNO) with the GNOT baseline and the Galerkin Transformer (G.T.). The left plot shows the evolution of relative L2 error over time steps for the SW2D equation, while the right plot shows the same for the SW2D equation. MNO consistently outperforms GNOT and G.T., particularly in terms of lower error growth, highlighting its superior ability to capture long-range dependencies and model complex dynamics more effectively.
Figure 10: Ablation studies on d_stated\_state(the hidden SSM state dimension), ssm_ratiossm\_ratio(the fraction of channels allocated to the SSM branch), and scan direction, showing trade-offs between RMSE (blue) and FLOPs (green).
@@ 4.4 Ablation Study: Final Battles, Winner Takes All
Table 5 compares nRMSE performance across two query scenarios: Identical positions (input and query points are the same) and Diagonal positions (shifted inputs creating a mismatch). Experiments are done using Darcy Flow. While prior work shows that Transformers handle inconsistent input-query positions well[li2022transformer], our results demonstrate a clear advantage of Mamba in both configurations. For Identical query positions, Mamba version achieves the lowest error, outperforming OFormer and its softmax variant, demonstrating Mamba’s superior ability to capture relationships when input and query points are perfectly aligned. For Diagonal query positions, where inputs and queries are misaligned, Mamba achieves the best performance compared to OFormer and its variant, demonstrating its superior ability to generalise under spatial shifts.
@@
Table 4 compares nRMSE performance across different dataset sizes for GNOT, GNOT with softmax attention (w/ S.A.), and GNOT with Mamba. Experiments are carried out using Darcy Flow. As dataset size decreases, Mamba consistently achieves the lowest error, demonstrating superior performance and robustness in data-scarce scenarios. For instance, with the smallest dataset (1K), Mamba achieves an nRMSE of 0.0617, significantly lower than GNOT’s and GNOT w/ S.A.’s, showcasing its resilience and generalisation capability even with limited data. This highlights Mamba’s efficiency in learning meaningful representations with fewer data points, making it a powerful choice for real-world applications where data availability is a constraint.
Computational Efficiency.
Across the three Transformer-based architectures—GNOT, Galerkin Transformer (G.T.), and OFormer—replacing the original attention mechanism with our Mamba Neural Operator (MNO) yields substantial efficiency gains. Compared to the softmax attention variants, MNO consistently reduces FLOPs, inference time, and GPU memory usage, often by an order of magnitude. For example, in OFormer, MNO lowers FLOPs from 977.14 G to just 60.13 G, reduces memory usage from 4.83 GiB to 1.13 GiB, and halves inference time. Similar trends are observed for GNOT and G.T., where MNO achieves up to a 10× reduction in FLOPs and over a 90 % decrease in memory footprint. When compared with the normalized attention method, MNO delivers approximately 50 % reductions in FLOPs, parameter count, and inference time, while maintaining similar GPU memory usage. Relative to the original Galerkin method, MNO exhibits higher FLOPs, a larger parameter count, and increases in both inference time and GPU memory usage. Theoretically, both Mamba and Galerkin methods have an O(N) memory cost; however, Mamba employs a bi-directional scan to enhance representational capacity, which in turn increases computational complexity relative to the Galerkin method. Despite this, Mamba consistently outperforms Galerkin in accuracy. Overall, these results demonstrate MNO’s ability to provide high‐accuracy PDE solutions with markedly improved computational efficiency, making it particularly well‐suited for large‐scale and real‐time applications.
Long-time stability.
The long-time stability of the PDE solvers is clearly improved by integrating our Mamba Neural Operator (MNO) into both GNOT and the Galerkin Transformer architectures. In our method, we adopts the one-shot prediction strategy, where the model predicts the entire future sequence directly from the initial condition:
$\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{T}\}=f_{\theta}(x_{0}).$
When training a one-shot prediction model, the network learns to capture correlations across the entire temporal sequence rather than progressing step-by-step. Because all future states are predicted jointly from the initial condition, the model can allocate its representational capacity to minimise the average error over the entire trajectory. In other words, it will not accumulate the error in each time step. This joint optimisation may produce a characteristic pattern where the error is slightly higher at early times (reflecting transient or complex dynamics) and then decreases as the prediction converges toward smoother, easier-to-represent states. Conceptually, this process resembles performing a global least-squares fit across all time steps simultaneously, rather than executing a sequential simulation where errors accumulate step-by-step. In the standard GNOT and Galerkin Transformer, the relative L2 error exhibits a noticeable increase during early time steps and remains at a higher baseline throughout the prediction horizon. In contrast, the MNO(GNOT+MNO and GalerkinTransformer+MNO) show a substantial reduction in error magnitude across all time steps, along with smaller error fluctuations. This reduced variance and lower error plateau indicate that the MNO-enhanced models maintain more accurate and stable predictions over extended time horizons, effectively mitigating error accumulation typical in autoregressive PDE forecasting. The effect is particularly pronounced in the Galerkin Transformer case, where MNO integration achieves over a two-fold reduction in long-term error levels.
Ablation Studies on Hyperparameter
The ablation study in Figure 10 evaluates the impact of three key hyperparameters - d_stated\_state, ssm_ratiossm\_ratio and scan direction - on model performance (RMSE) and computational cost (FLOP). d_stated\_state means the dimensionality of the latent hidden state in the state-space model (SSM) branch of Mamba. A larger d_stated\_state means that the SSM maintains a richer internal memory representation at each step, potentially capturing more complex temporal dependencies, but at the cost of higher computational complexity. ssm_ratiossm\_ratio means the proportion of the model’s total feature channels ($d_{\text{model}}$) that are allocated to the SSM branch within each Mamba block. For d_stated\_state, increasing the hidden SSM state dimension from 12 to 24 leads to higher FLOPs and a non-monotonic effect on RMSE, with the best accuracy achieved at $d_{state}=16$. Similarly, raising the ssm_ratiossm\_ratio from 1 to 4 steadily increases FLOPs, with the lowest RMSE observed at ssm_ratio=2ssm\_ratio=2, beyond which accuracy degrades. Regarding scan direction, the four-directional variant results in slightly higher RMSE and increased FLOPs compared to the bi-directional scan, indicating that the added complexity does not translate to better accuracy in this setup. Overall, the results suggest that our model is the best choice.
4.5 Limitation
Irregular Domain.
[…]
5 Conclusion
@@ Acknowledgments
CWC and AIAR acknowledge support from the Swiss National Science Foundation (SNSF) under grant number 20HW-1 220785. JH was supported in part by the Imperial College Bioengineering Department PhD Scholarship and the UKRI Future Leaders Fellowship(MR/V023799/1). GY was supported in part by the ERC IMI (101005122), the H2020 (952172), the MRC (MC/PC/21013), theRoyal Society (IEC%NSFC%211235), the NVIDIA Academic Hardware Grant Program, the SABER project supported by BoehringerIngelheim Ltd, NIHR Imperial Biomedical Research Centre (RDA01), The Wellcome Leap Dynamic resilience program (co-funded by Temasek Trust), UKRI guarantee funding for Horizon Europe MSCA Postdoctoral Fellowships (EP/Z002206/1), UKRI MRC ResearchGrant, TFS Research Grants (MR/U506710/1), Swiss National Science Foundation (Grant No. 220785), and the UKRI Future Leaders Fellowship (MR/V023799/1, UKRI2738). CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC advanced career fellowship EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards 215733/Z/19/Z and 221633/Z/20/Z, CCMI, and the Alan Turing Institute. AIAR gratefully acknowledges the support of the Yau Mathematical Sciences Center, Tsinghua University. This work is also supported by the Tsinghua University Dushi Program.
Appendix A. Notation Table
To improve readability and address any ambiguity, we summarise in Table 7 the notation used throughout the paper. The table groups symbols by context (Mamba Neural Operator, General PDE and hyperparameters) and provides a concise description of their meaning. This helps ensure consistency across sections and serves as a quick reference for the reader.
Appendix B. More explanation on the main method
In this section, we provides an additional example on how the difference between Euler and ZOH method. Consider the scalar linear state–space model
${h}^{\prime}(t)=ah(t)+bu(t),$ (34)
with a=1a=-1, $b=2$. We look at one step of size $\Delta>0$ from $t_{k}$ to $t_{k+1}=t_{k}+\Delta$. For the Forward Euler, we defines as follows:
$h_{k+1}^{\text{Euler}}=(1+a\Delta)\,h_{k}+b\Delta\,u_{k}.$ (35)
For the Zero-Order Hold (ZOH), we defines as follows:
$h_{k+1}^{\text{ZOH}}=e^{a\Delta}\,h_{k}+\left(\int_{0}^{\Delta}e^{a\tau}\,d\tau\right)bu_{k}=e^{a\Delta}\,h_{k}+\frac{e^{a\Delta}-1}{a}\,bu_{k},$ (36)
which corresponds to the matrix pair $(\tilde{A},\tilde{B})=\big(e^{A\Delta},\;A^{-1}(e^{A\Delta}-I)B\big)$ in the scalar case.
If we truncate the exponential at first order,
$e^{a\Delta}\approx 1+a\Delta,\qquad\frac{e^{a\Delta}-1}{a}\approx\Delta,$ (37)
we recover the Euler pair $(I+\Delta A,\,\Delta B)$.
If we Let $u(t)\equiv 1$ on $t_{k},t_{k+1})$ with arbitrary $h_{k}$.
$h_{k+1}^{\text{ZOH}}$ $=e^{-\Delta}h_{k}+2(1-e^{-\Delta}),$
$h_{k+1}^{\text{Euler}}$ $=(1-\Delta)h_{k}+2\Delta.$
Then Using Talyr expansion, we obtain
$\underbrace{h_{k+1}^{\text{ZOH}}}_{\text{exact}}=\Big(1-\Delta+\tfrac{\Delta^{2}}{2}-\cdots\Big)h_{k}+2\Big(\Delta-\tfrac{\Delta^{2}}{2}+\tfrac{\Delta^{3}}{6}-\cdots\Big).$ (38)
Euler keeps only the first-order pieces: $h_{k+1}^{\text{Euler}}=(1-\Delta)h_{k}+2\Delta$. Then the per-step difference is $h_{k+1}^{\text{ZOH}}-h_{k+1}^{\text{Euler}}=\tfrac{\Delta^{2}}{2}\,h_{k}-\Delta^{2}+O(\Delta^{3})=O(\Delta^{2})$.
Numerical illustration (from $h_{k}=0$).
  • $\Delta=0.5$:  $h_{k+1}^{\text{ZOH}}=2(1-e^{-0.5})\approx 0.78694$,   $h_{k+1}^{\text{Euler}}=1.0$,   error $\approx 0.21306$.
  • $\Delta=0.25$:  $h_{k+1}^{\text{ZOH}}\approx 0.44240$,   $h_{k+1}^{\text{Euler}}=0.5$,   error $\approx 0.05760$.
As $\Delta$ halves, the error shrinks by a factor of $\approx 4$ (consistent with $O(\Delta^{2})$ local error), while ZOH is exact under the constant-input assumption.
Appendix C. More Time Analysis Results
Consistent with the main section, we provide additional long-time stability analysis on a different dataset with an alternative baseline. Figure 11 illustrates the long-time integration stability of OFormer and Galerkin Transformer, both with and without the incorporation of our proposed MNO. The OFormer baseline demonstrates a noticeably faster convergence rate during training compared to our proposed model. This behavior can be attributed to OFormer’s normalization-based operator formulation, which effectively stabilizes the optimization dynamics by rescaling both the input and operator spectra. Such normalization reduces gradient variance and facilitates smoother propagation through layers, thereby accelerating convergence. Although OFormer exhibits a steeper early decrease in relative L2 error after the initial spike (faster short‑horizon convergence), our proposed Mamba neural operator suppress that spike and maintain lower error growth over the rollout. For the OFormer backbone the asymptotic errors are comparable, but MNO avoids the large early overshoot and yields a smoother, more stable trajectory. In both cases, the baseline models (blue curves) exhibit rapid error growth during extended rollout, with the error plateauing at a relatively high level or even drifting upward. By contrast, when leveraging Mamba neural operator (orange curves), the models maintain substantially lower error throughout the entire temporal horizon. The error remains bounded and exhibits minimal drift, indicating that MNO effectively mitigates the accumulation of integration errors over long rollouts. This demonstrates that our method not only improves short-term predictive accuracy but also provides enhanced numerical stability and robustness for long-time forecasting of dynamical systems.
Table 7: Comprehensive notation reference for symbols used throughout the manuscript. The table provides quick lookup for mathematical definitions and aligns notation across different sections.
Notation Description
$h(t)\in\mathbb{R}^{N}$ State vector at continuous time $t$
$u(t)\in\mathbb{R}^{L}$ Control/input vector
$y(t)\in\mathbb{R}^{L}$ Output vector
$A\in\mathbb{C}^{N\times N}$ State transition matrix
$B\in\mathbb{C}^{N}$ Input-to-state matrix
$C\in\mathbb{C}^{N}$ State-to-output matrix
$D\in\mathbb{C}^{N}$ Direct input-to-output matrix
$\Delta$ Time discretisation step size
$x_{k}\approx x(t_{k})$ State at discrete time index $k$, with $t_{k}=k\Delta$
$u_{k}\approx u(t_{k})$ Input at discrete time index $k$
$y_{k}\approx y(t_{k})$ Output at discrete time index $k$
$A_{d}=e^{A\Delta}$ Discretised state transition matrix (from matrix exponential)
$B_{d}=\int_{0}^{\Delta}e^{A\tau}B\,d\tau$ Discretised input matrix
$x_{k+1}=A_{d}x_{k}+B_{d}u_{k}$ Discrete-time state update equation
$y_{k}=Cx_{k}+Du_{k}$ Discrete-time observation equation
Hyperparameters
d_stated\_state size of the SSM hidden state
ssm_ratiossm\_ratio share of $d_{\text{model}}$ channels in the SSM branch
Figure 11: Relative L2 error on DR2D across time steps, comparing the performance of our Mamba Neural Operator (MNO) with the OFormer baseline and the Galerkin Transformer (G.T.). The left plot shows the evolution of relative L2 error over time steps for the DR2D equation, while the right plot shows the same for the DR2D equation. MNO consistently outperforms OFormer and G.T., particularly in terms of lower error growth, highlighting its superior ability to capture long-range dependencies and model complex dynamics more effectively.